15y - 11 = 10y - 7 ?
Found 2 solutions by stanbon, mxgirl22:
Answer by stanbon(75887)

15y - 11 = 10y - 7
----
5y = 4
----
y = 4/5
---------------
Cheers,
Stan H.
--------------

Answer by mxgirl22(39)

15y - 11 = 10y - 7 ?

------------------------------------------------

Step 1. Get the y to one side of the equal sign


15y - 11 = 10y - 7

15y - 11 -10y = 10y - 10y - 7

5y - 11 = -7
-----------------------------------------------
Step 2. Get the y by itself.

5y -11 +11 = -7 + 11

5y = 4
-----------------------------------------------

Step 3. Solve for y.

5y/5 = 4/5

y = 4/5
-----------------------------------------------

Step 4. Answer

y = 4/5
----------------------------------------------
Step 5. Check with original problem


15y - 11 = 10y - 7


15(4/5) - 11 =10(4/5) - 7


12 - 11 =10(4/5) - 7


1 = 8 - 7


1 = 1


Answer checks!